OpenStudy (anonymous):

Given two spheres, if the volume of the first sphere is 4/3*3.14 and the volume of the second sphere is 36*3.14 , which is the relation of the second radius to the first? squared tripled doubled multiplied by a factor of 8/3

OpenStudy (anonymous):

\[\frac{4}{3}pi r^3 = \frac{4}{3}\pi\] So for this first sphere, r must be equal to 1. \[\frac{4}{3}\pi r^3 = 36\pi\] You need to solve this second equation for r. Cancel out the pi and Multiply both sides by (3/4) to get \[r^3 = 27\] \[r = \sqrt[3]{27} = \sqrt[3]{3^3} = 3\] r = 3 for the second sphere. The second radius is triple the first radius.